OpenStudy (reverdeparadis):

0.5(m + 4) = 3(m - 1) a.) -2 b.) 2 c.) no solution d.) identity

OpenStudy (anonymous):

Solve\[\frac{5}{10}(m+4)=3(m-1) \]for m.

OpenStudy (anonymous):